OpenStudy (anonymous):

what is the equation of a line that passes through the point (2,1) and ahs a slope of 2?

hartnn (hartnn):

the slope point form of line is y-y1 =m (x-x1) here m =slope =2 x1 = 2, y1 = 1 just plug in values! what do u get ?

hartnn (hartnn):

and bdw, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)